a vending machine has 12oz drinks and 16oz drinks. The 12oz drinks cost $.50 cents and the 16oz drinks cost 1 dollar. If 3600 drinks were sold in one month and total sales were $2225, how many of each size drink was sold?

Answer by checkley79(3341)

X+Y=3,600
X=3,600-Y
.50X+1Y=2,225
.50(3,600-Y)+Y=2,225
1,800-.50Y+Y=2,225
.50Y=2,225-1,800
.50Y=425
Y=425/.5
Y=850 $1.00 BOTTLES WERE SOLD.
X=3,600-850=2,750 $.50 BOTTLES WERE SOLD.
PROOF:
.50*2,750+1*850
1,375+850=2,225
2,225=2,225